# Convergence of Stochastic Gradient Langevin Dynamics in the Lazy Training Regime

### v1

@@ 1.2 Notation
IdId denotes the identity matrix. $A\preceq B$ indicates that BAB-A is positive-definite. $\|x\|_{2}=\sqrt{x^{T}x}$ denotes the 2-norm of $x\in\mathbb{R}$. We use $\lambda_{min}(A)$ for the smallest eigenvalue of a Matrix $A$. For a positive-definite matrix $A$, we define the vector norm $\|x\|_{A}^{2}:=x^{T}Ax$. Additionally $\|A\|_{F}$ is the Frobenius norm of $A$ and $B_{r}(x_{0})$ is the open ball of radius $r$ around the $x_{0}$, in the Euclidean norm. We denote the modulus of Lipschitz continuity of a function $f$ as $\mathrm{Lip}(f)$. Lastly, $\mathds{1}_{A}$ denotes the characteristic function, which is equal to $1$ if the logical expression $A$ holds true and 0 otherwise.
@@ 2.1 Problem Setting
$\omega^{\star}=\operatorname*{arg\,min}_{\omega\in\mathbb{R}^{p}}R(h(\omega)).$
Empirical risk minimization in supervised learning is a special case of the formulation above, where $P$ is a discrete probability measure on a given set of data points $\{(x_{i},y_{i})\}_{i=1}^{n}$ and $h(\omega^{\star})(x_{i})=y_{i}$. For further details about the supervised learning formulation, see Section 4, where we apply the results to train neural networks with smooth activations in a supervised learning setting.
@@ 2.2 Lazy Training as Stochastic Langevin Dynamics
$d\omega_{t}=-\frac{1}{\alpha}D^{T}h(\omega_{t})\nabla R(\alpha h(\omega_{t}))dt+\frac{\sqrt{\eta_{\alpha}}}{\alpha}(\Sigma_{\alpha}(\omega_{t}))^{\frac{1}{2}}dW_{t},$ (1)
where $W_{t}$ is a standard Brownian motion in $\mathbb{R}^{p}$, $\Sigma_{\alpha}(\omega)$ is the covariance matrix of $\nabla\ell(\cdot,\alpha h(\omega))$ and $\eta_{\alpha},\alpha\in\mathbb{R}_{+}$. We adopt a random initialization $\omega_{0}$ such that $h(\omega_{0})=0$ holds almost surely. For neural networks, this can be achieved by using the symmetric initialization scheme (Bai & Lee, 2020; Chizat et al., 2020; Cayci & Eryilmaz, 2025).
Here, $\sigma_{\alpha}(\omega):=(\Sigma_{\alpha}(\omega))^{\frac{1}{2}}$ denotes the matrix square root $\sigma_{\alpha}^{T}(\omega)\sigma_{\alpha}(\omega)=\Sigma_{\alpha}(\omega)$, which is well-defined since $\Sigma_{\alpha}(\omega)$ is symmetric positive definite. The noise scaling factor $\eta_{\alpha}$ may depend on the output scaling factor $\alpha$ in certain cases (see Appendix 6.1).
@@
$\omega_{k+1}=\omega_{k}-\frac{\eta_{\alpha}}{\alpha}D^{T}h(\omega_{k})\nabla R(\alpha h(\omega_{k}))+\frac{\sqrt{\eta_{\alpha}}}{\alpha}V_{k},$ (2)
where $V_{k}:=\sqrt{\eta_{\alpha}}(D^{T}h(\omega_{k})\nabla R(\alpha h(\omega_{k}))-D^{T}h(\omega_{k})\nabla l(x,\alpha h(\omega_{k})))$ and $x\sim\mathrm{Unif}\{x_{1},\ldots,x_{n}\}$ is a sample from the distribution on the input space. In that case, $V_{k}|x_{k}$ is a random vector with mean 0 and covariance matrix $\eta_{\alpha}\Sigma_{\alpha}(x_{k})$. In Li et al. (2017) it is shown that the stochastic continuous-time model that we consider in equation 2 is an order-one weak approximation of the SDE equation 1. There are other stochastic models to approximate SGD in continuous time (Li et al., 2017; Simsekli et al., 2019; Gurbuzbalaban et al., 2021).
@@ 3 Convergence of Stochastic Gradient Langevin Dynamics in the Lazy Training Regime
    $\nabla\ell(x,h_{1})-\nabla\ell(x,h_{2})||_{2}<\mathrm{Lip}(\nabla\ell)||h_{1}-h_{2}||_{\mathbb{F}},$
@@
Assumption 1 holds for many loss functions, with the most common example being the squared error (see Appendix 6.1). The smoothness assumption on $h$ holds for shallow neural networks (see Corollary 1). For deep neural networks, we will present an alternative that circumvents the smoothness of $h$ (see Corollary 2 and Appendix 6.2). The empirical neural tangent kernel is strictly positive definite for overparameterized neural networks with high probability over the random initialization under reasonable data regularity conditions (Du et al., 2019b; a; Banerjee et al., 2023).
@@
We assume that $\nabla^{2}_{\omega}R(\alpha h(\omega))\preceq\frac{\alpha^{2}}{\eta_{\alpha}}Id$ for all $\omega\in B_{r}(\omega_{0})$, where $\lambda^{2}=\lambda_{min}(Dh^{T}(\omega_{0})Dh(\omega_{0}))$ and $r:=\frac{\lambda}{\mathrm{Lip}(Dh)}.$
It can easily be shown that Assumption 2 holds for shallow and deep neural networks. In Appendix 6.1, we prove that Assumption 2 holds for shallow neural networks with smooth activation functions even outside of the ball $B_{r}(\omega_{0})$.
@@
Assume that the gradient and the expectation of the loss interchange, i.e., $\nabla R(h)=\mathbb{E}[\nabla\ell(x,h)]$.
@@
Under Assumption 1 and Assumption 3, $R$ is $\mathrm{Lip}(\nabla\ell)$-smooth and m-strongly convex.
@@
$\|\nabla R(h_{1})-\nabla R(h_{2})\|$ $=\|\mathbb{E}_{x}[\nabla\ell(x,h_{1})-\nabla\ell(x,h_{2})]\|_{2}$
$\leq\mathbb{E}[\|\nabla\ell(x,h)-\nabla\ell(x,h_{2}))\|_{2}]$
$\leq\mathbb{E}_{x}[\mathrm{Lip}(\nabla\ell)\|h_{1}-h_{2}\|_{\mathbb{F}}]$
$=\mathrm{Lip}(\nabla\ell)\|h_{1}-h_{2}\|_{\mathbb{F}}.$
@@
$(\nabla R(h_{1})-\nabla R(h_{2}))^{T}(h_{1}-h_{2})$ $=\mathbb{E}_{x}[(\nabla\ell(x,h_{1})-\nabla\ell(x,h_{2}))^{T}](h_{1}-h_{2})$
$=\mathbb{E}_{x}[(\nabla\ell(x,h_{1})-\nabla\ell(x,h_{2}))^{T}(h_{1}-h_{2})]$
@@
As a key feature of analysis in the lazy training regime, the weights $\omega_{t}$ do not stray too far from their initial value, which implies that $\lambda Id\preceq Dh(\omega_{t})Dh^{T}(\omega_{t})$ for some $\lambda>0$. Many publications have focused on establishing strictly positive lower bounds for this value for deterministic (full-batch) gradient flow (see, for instance, Du et al. (2019b); Chizat et al. (2020)); however, we are not aware of any results that guarantee the positivity of $\lambda=\lambda_{min}(Dh(\omega_{t})Dh^{T}(\omega_{t}))$ for SGD or SGLD. We will solve this problem by introducing a stopping time, which is the smallest time at which this condition no longer holds, and prove in Corollary 3 that this stopping time will be $\infty$ with high probability for $\alpha$ sufficiently large. The minimum eigenvalue of the NTK is strictly positive, at least as long as $\omega_{t}-\omega_{0}||\leq r:=\frac{\lambda}{\mathrm{Lip}(Dh)}$. Therefore, we will introduce the random stopping time
@@
denoting the first time $\lambda Id\preceq Dh(\omega_{t})D^{T}h(\omega_{t})$ is degenerate. In addition, $\nabla^{2}_{\omega}R(\alpha h(\omega_{t}))\preceq\frac{\alpha^{2}}{\eta_{\alpha}}Id$ also holds at least for $t\leq\tau$ under Assumption 2. Working with this quantity and proving $\tau=\infty$ with high probability is a major challenge in our analysis. In addition, we define $\min\{T,\tau\}=:\tilde{T}$, where $T>0$ is some arbitrary but fixed runtime of the algorithm. Let $t<\tilde{T}$ in the following.
@@
$\mathbb{E}[\bar{R}(\alpha h(\omega_{t}))]\leq\mathbb{E}[\bar{R}(\alpha h(\omega_{0}))]\exp(-2m\lambda^{2}t),$ (3)
@@
As defined, $\omega_{t}$ follows the SDE in equation 1. In the following, we apply Itô’s formula to compute $dg(\omega_{t})$, where $g:\mathbb{R}^{p}\rightarrow\mathbb{R}$ is defined by $g(\omega)=\log(\bar{R}(\alpha h(\omega)))$. The gradient and the Hessian of $g$ are as follows:
$\nabla g(\omega)=$ $\frac{1}{\bar{R}(\alpha h(\omega))}\alpha D^{T}h(\omega)\nabla R(\alpha h(\omega))$
$\nabla^{2}_{\omega}g(\omega)=$ $-\frac{1}{\bar{R}(\alpha h(\omega))^{2}}\alpha^{2}D^{T}h(\omega)\nabla R(\alpha h(\omega))\nabla^{T}R(\alpha h(\omega))Dh(\omega)+\frac{1}{\bar{R}(\alpha h(\omega))}\nabla^{2}_{\omega}R(\alpha h(\omega))$
@@
$dg(\omega_{t})=$ $-\frac{\|D^{T}h(\omega_{t})\nabla R(\alpha h(\omega_{t}))\|^{2}}{\bar{R}(\alpha h(\omega_{t}))}dt$ (4)
$+\frac{\eta_{\alpha}}{2}\mathrm{Tr}\left(-\frac{\|\nabla^{T}R(\alpha h(\omega_{t}))Dh(\omega_{t})\sigma_{\alpha}(\omega_{t})\|_{2}^{2}}{\bar{R}(\alpha h(\omega_{t}))^{2}}\right)dt$ (5)
@@
$+\frac{\sqrt{\eta_{\alpha}}\nabla^{T}R(\alpha h(\omega_{t}))Dh(\omega_{t})\sigma_{\alpha}(\omega_{t})}{\bar{R}(\alpha h(\omega_{t}))}dW_{t}.$ (7)
@@
$M_{t}:=\sqrt{\eta_{\alpha}}\int_{0}^{t}\frac{\nabla^{T}R(\alpha h(\omega_{s}))Dh(\omega_{t})\sigma_{\alpha}(\omega_{s})}{\bar{R}(\alpha h(\omega_{s}))}dW_{s},$
@@
$\langle M\rangle_{t}=\eta_{\alpha}\int_{0}^{t}\mathrm{Tr}\left(-\frac{\|\nabla^{T}R(\alpha h(\omega_{s}))Dh(\omega_{s})\sigma_{\alpha}(\omega_{s}))\|_{2}^{2}}{\bar{R}(\alpha h(\omega_{s}))^{2}}\right)ds,$
@@
$\bar{R}(\alpha h(\omega_{t}))=$ $\bar{R}(\alpha h(\omega_{0}))\exp{\left(-\int_{0}^{t}\frac{\|D^{T}h(\omega_{s})\nabla R(\alpha h(\omega_{s}))\|_{2}^{2}}{\bar{R}(\alpha h(\omega_{s}))}ds\right)}$
@@
$\int_{0}^{t}\frac{\nabla^{T}R(\alpha h(\omega_{s}))Dh(\omega_{s})D^{T}h(\omega_{s})\nabla R(\alpha h(\omega_{s}))}{\bar{R}(\alpha h(\omega_{s}))}ds=\int_{0}^{t}\frac{\|\nabla R(\alpha h(\omega_{s}))\|^{2}_{Dh(\omega_{s})Dh^{T}(\omega_{s})}}{\bar{R}(\alpha h(\omega_{s}))}ds.$
@@
$\int_{0}^{t}\frac{\|\nabla R(\alpha h(\omega_{s}))\|^{2}_{Dh(\omega_{s})Dh(\omega_{s})^{T}}}{\bar{R}(\alpha h(\omega_{s}))}ds\geq\lambda^{2}\int_{0}^{t}\frac{\|\nabla R(\alpha h(\omega_{s}))\|^{2}_{2}}{\bar{R}(\alpha h(\omega_{s}))}\geq 2\lambda^{2}m\int_{0}^{t}ds=2\lambda^{2}mt,$
using the Polyak-Lojasiewicz inequality, which holds due to the strong convexity of $R$. Next we bound
$\frac{\eta_{\alpha}}{2\alpha^{2}}\int_{0}^{t}\mathrm{Tr}\left(\frac{\sigma_{\alpha}(\omega_{s})^{T}\nabla^{2}_{\omega}R(\alpha h(\omega_{s}))\sigma_{\alpha}(\omega_{s})}{\bar{R}(\alpha h(\omega_{s}))}\right)ds=$ $\frac{\eta_{\alpha}}{2\alpha^{2}}\int_{0}^{t}\mathrm{Tr}\left(\nabla^{2}_{\omega}R(\alpha h(\omega_{s}))\frac{\Sigma_{\alpha}(\omega_{s})}{\bar{R}(\alpha h(\omega_{s}))}\right)ds$
$\leq$ $\frac{\eta_{\alpha}}{2\alpha^{2}}\int_{0}^{t}\mathrm{Tr}\left(\nabla^{2}_{\omega}R(\alpha h(\omega_{s}))\right)\mathrm{Tr}\left(\frac{\Sigma_{\alpha}(\omega_{s})}{\bar{R}(\alpha h(\omega_{s}))}\right)ds.$
By using that $\nabla^{2}_{\omega}R(\alpha h(\omega_{t}))\preceq\frac{\alpha^{2}}{\eta_{\alpha}}Id$ and using the definition of $\Sigma_{\alpha}(\omega)$, we get
$\frac{\eta_{\alpha}}{2\alpha^{2}}\int_{0}^{t}\mathrm{Tr}\left(\nabla^{2}_{\omega}R(\alpha h(\omega_{s}))\right)$ $\mathrm{Tr}\left(\frac{\Sigma_{\alpha}(\omega_{s})}{\bar{R}(\alpha h(\omega_{s}))}\right)ds$
$\leq$ $\frac{1}{2}\int_{0}^{t}\mathrm{Tr}\left(\frac{\Sigma_{\alpha}(\omega_{s})}{\bar{R}(\alpha h(\omega_{s}))}\right)ds$
$=$ $\frac{1}{2}\int_{0}^{t}\frac{\mathrm{Tr}\left(\mathbb{E}_{x}[\nabla\ell(x,\alpha h(\omega_{s}))\nabla^{T}\ell(x,\alpha h(\omega_{s}))]-\nabla R(\alpha h(\omega_{s}))\nabla^{T}R(\alpha h(\omega_{s}))\right)}{\bar{R}(\alpha h(\omega_{s}))}ds$
$=$ $\frac{1}{2}\int_{0}^{t}\frac{\mathbb{E}_{x}[\|\nabla\ell(x,\alpha h(\omega_{s}))\|^{2}_{2}]-\|\nabla R(\alpha h(\omega_{s}))\|^{2}_{2}]}{\bar{R}(\alpha h(\omega_{s}))}ds$
$\overset{(\clubsuit)}{\leq}$ $\frac{1}{2}\int_{0}^{t}\frac{\mathbb{E}_{x}[2m(\ell(x,\alpha h(\omega_{s}))-\ell^{\star}(x))]}{\mathbb{E}_{x}[\ell(x,\alpha h(\omega_{s}))-\ell^{\star}(x)]}-2m\ ds$
$=$ $0,$
where the inequality in $(\clubsuit)$ follows from the Polyak-Łojasiewicz inequality on the $m$-strongly convex function $\ell$. Consequently we obtain
$\bar{R}(\alpha h(\omega_{t}))\leq\bar{R}(\alpha h(\omega_{0}))\exp{\left(-2\lambda^{2}mt+M_{t}-\frac{1}{2}\langle M\rangle_{t}\right)}.$
@@
$\mathbb{E}[\bar{R}(\alpha h(\omega_{t})]\xrightarrow[]{t\rightarrow\infty}0,$

### v2

@@ 1.1 Related Work
Our work is closely related to Chizat et al. (2020) and Lugosi & Nualart (2024). However, there are many substantial differences between prior work and our contributions. For instance, Chizat et al. (2020) does not consider the stochastic term needed to model stochastic learning algorithms. Compared to Lugosi & Nualart (2024), we investigate the dynamics in the lazy training regime, which allows us to formulate less restrictive and easily verifiable assumptions in a machine learning context. An indepth discussion on the differences of our contributions and prior works is presented in Appendix A.
@@ 1.2 Notation
IdId denotes the identity matrix. $A\preceq B$ indicates that BAB-A is positive-definite. $\|x\|_{2}=\sqrt{x^{T}x}$ denotes the 2-norm of $x\in\mathbb{R}$. We use $\lambda_{min}(A)$ for the smallest eigenvalue of a Matrix $A$. For a positive-definite matrix $A$, we define the vector norm $\|x\|_{A}^{2}:=x^{T}Ax$. Additionally, $\|A\|_{F}$ is the Frobenius norm of $A$ and $B_{r}(x_{0})$ is the open ball of radius $r$ around the $x_{0}$, in the Euclidean norm. We denote the modulus of Lipschitz continuity of a function $f$ as $\mathrm{Lip}(f)$. The notation $Dh$ denotes the Jacobian of $h$. Similarly, $\nabla h$ represents the gradient of $h$ and $\nabla^{2}h$ represents its Hessian. For a function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$, it holds that $\nabla^{T}f=Df$. We add an index to $\nabla$ to indicate, that the gradient is only taken with respect to one of the arguments. $\nabla_{h}$ can be seen as the gradient in the Hilbert space $\mathbb{F}$. However, this can be replaced by the standard gradient with respect to the values that $h\in\mathbb{F}$ takes on the finite set $\{x_{i}\}_{i=1}^{n}$. Lastly, $\mathds{1}_{A}$ denotes the characteristic function, which is equal to $1$ if the logical expression $A$ holds true and 0 otherwise.
@@ 2.1 Problem Setting
$\mathrm{Find}\ \omega^{\star}\ \mathrm{s.t.}\ R(h(\omega^{\star}))<\epsilon,$
for some $\epsilon>0$. Empirical risk minimization in supervised learning is a special case of the formulation above, where $P$ is a discrete probability measure on a given set of data points $\{(x_{i},y_{i})\}_{i=1}^{n}$ and an unknown ground truth $h^{\star}\in\mathbb{F}$, that connects $x_{i}$ and $y_{i}$ by $h^{\star}(x_{i})=y_{i}$. In empirical risk minimization, a popular choice for the loss function is given by $\ell(x,h)=(h(x)-y)^{2}$, where $y=h^{\star}(x)$. For further details about the supervised learning formulation, see Section 4, where we apply the results to train neural networks with smooth activations in a supervised learning setting.
@@ 2.2 Lazy Training as Stochastic Langevin Dynamics
$d\omega_{t}=-\frac{1}{\alpha}D^{T}h(\omega_{t})\nabla_{h}R(\alpha h(\omega_{t}))dt+\frac{\sqrt{\eta_{\alpha}}}{\alpha}(\Sigma_{\alpha}(\omega_{t}))^{\frac{1}{2}}dW_{t},$ (1)
where $W_{t}$ is a standard Brownian motion in $\mathbb{R}^{p}$, $\Sigma_{\alpha}(\omega)$ is the covariance matrix of $\nabla_{h}\ell(\cdot,\alpha h(\omega))$ and $\eta_{\alpha},\alpha\in\mathbb{R}_{+}$. We adopt a random initialization $\omega_{0}$ such that $h(\omega_{0})=0$ holds almost surely. For neural networks, this can be achieved by using the symmetric initialization scheme (Bai & Lee, 2020; Chizat et al., 2020; Cayci & Eryilmaz, 2025).
Here, $\sigma_{\alpha}(\omega):=(\Sigma_{\alpha}(\omega))^{\frac{1}{2}}$ denotes the matrix square root $\sigma_{\alpha}^{T}(\omega)\sigma_{\alpha}(\omega)=\Sigma_{\alpha}(\omega)$, which is well-defined since $\Sigma_{\alpha}(\omega)$ is symmetric positive definite. The noise scaling factor $\eta_{\alpha}$ may depend on the output scaling factor $\alpha$ in certain cases (see Appendix B).
@@
$\omega_{k+1}=\omega_{k}-\frac{\eta_{\alpha}}{\alpha}D^{T}h(\omega_{k})\nabla_{h}R(\alpha h(\omega_{k}))+\frac{\sqrt{\eta_{\alpha}}}{\alpha}V_{k},$ (2)
where $V_{k}:=\sqrt{\eta_{\alpha}}(D^{T}h(\omega_{k})\nabla_{h}R(\alpha h(\omega_{k}))-D^{T}h(\omega_{k})\nabla_{h}l(x,\alpha h(\omega_{k})))$ and $x\sim\mathrm{Unif}\{x_{1},\ldots,x_{n}\}$ is a sample from the distribution on the input space. In that case, $V_{k}|x_{k}$ is a random vector with mean 0 and covariance matrix $\eta_{\alpha}\Sigma_{\alpha}(x_{k})$. In Li et al. (2017) it is shown that the stochastic continuous-time model that we consider in equation 2 is an order-one weak approximation of the SDE equation 1. There are other stochastic models to approximate SGD in continuous time (Li et al., 2017; Simsekli et al., 2019; Gurbuzbalaban et al., 2021).
@@ 3 Convergence of Stochastic Gradient Langevin Dynamics in the Lazy Training Regime
    $\nabla_{h}\ell(x,h_{1})-\nabla_{h}\ell(x,h_{2})||_{2}<\mathrm{Lip}(\nabla_{h}\ell)||h_{1}-h_{2}||_{\mathbb{F}},$
@@
Assumption 1 holds for many loss functions, with the most common example being the squared error (see Appendix B). The smoothness assumption on $h$ holds for shallow neural networks (see Corollary 1). For deep neural networks, we will present an alternative that circumvents the smoothness of $h$ (see Corollary 2 and Appendix C). The empirical neural tangent kernel is strictly positive definite for overparameterized neural networks with high probability over the random initialization under reasonable data regularity conditions (Du et al., 2019b; a; Banerjee et al., 2023).
@@
We assume that $Tr\left(\frac{\sigma_{\alpha}(\omega)^{T}\nabla_{\omega}^{2}R(\alpha h(\omega))\sigma_{\alpha}(\omega)}{\bar{R}(\alpha h(\omega))}\right)\leq\frac{2\alpha^{2}\lambda^{2}m}{\eta_{\alpha}}$, for all $\omega\in B_{r}(\omega_{0})$, where $r$ is the radius that will be discussed later in this section (before Theorem 1).
Assumption 2 automatically holds for sufficiently over parametrized shallow and deep neural networks under realistic assumptions. In Appendix B and C, we provide the proof for shallow and deep neural networks with smooth activation functions.
@@
Assume that the gradient and the expectation of the loss interchange, i.e., $\nabla_{h}R(h)=\mathbb{E}[\nabla_{h}\ell(x,h)]$.
@@
Under Assumption 1 and Assumption 3, $R$ is $\mathrm{Lip}(\nabla_{h}\ell)$-smooth and m-strongly convex.
@@
$\|\nabla_{h}R(h_{1})-\nabla_{h}R(h_{2})\|$ $=\|\mathbb{E}_{x}[\nabla_{h}\ell(x,h_{1})-\nabla_{h}\ell(x,h_{2})]\|_{2}$
$\leq\mathbb{E}_{x}[\|\nabla_{h}\ell(x,h_{1})-\nabla_{h}\ell(x,h_{2}))\|_{2}]$
$\leq\mathbb{E}_{x}[\mathrm{Lip}(\nabla_{h}\ell)\|h_{1}-h_{2}\|_{\mathbb{F}}]$
$=\mathrm{Lip}(\nabla_{h}\ell)\|h_{1}-h_{2}\|_{\mathbb{F}}.$
@@
$(\nabla_{h}R(h_{1})-\nabla_{h}R(h_{2}))^{T}(h_{1}-h_{2})$ $=\mathbb{E}_{x}[(\nabla_{h}\ell(x,h_{1})-\nabla_{h}\ell(x,h_{2}))^{T}](h_{1}-h_{2})$
$=\mathbb{E}_{x}[(\nabla_{h}\ell(x,h_{1})-\nabla_{h}\ell(x,h_{2}))^{T}(h_{1}-h_{2})]$
@@
Remark 2.
At this point, it is essential to point out that Assumption 1 is already sufficient for the existence of a global minimizer of $R$ in the supervised learning setting with a finite set of data points. Since we have shown in Lemma 1 that $h\mapsto R(h)$ is $m$-strong convex as a function on the prediction space. Since the prediction space can be characterized by a final dimensional set if $R$ only depends on a finite number of values of $h$, we already know that there exists a global minimizer $h^{\star}\in\mathbb{F}$. However, we do not assume the existence of $\omega^{\star}\in\mathbb{R}^{p}$ such that $h(\omega^{\star})=h^{\star}$ (which is not needed for the theoretical findings of this paper). The existence of $\omega^{\star}$ is generally not guaranteed, since the mapping $\omega\mapsto R(h(\omega))$ does not have to be strongly convex. Later in the numerical experiments as well as the proofs of Corollaries 1 and 2, we assumed the existence of $\omega^{\star}$, since it holds for the teacher-student setting, and to simplify the calculations.
As a key feature of analysis in the lazy training regime, the weights $\omega_{t}$ do not deviate too far from their initial value, which implies that $\lambda Id\preceq Dh(\omega_{t})D^{T}h(\omega_{t})$ for some $\lambda>0$. Many publications have focused on establishing strictly positive lower bounds for this value for deterministic (full-batch) gradient flow (see, for instance, Du et al. (2019b); Chizat et al. (2020)); however, we are not aware of any results that guarantee the positivity of $\lambda=\lambda_{min}(Dh(\omega_{t})D^{T}h(\omega_{t}))$ for SGD or SGLD. We will solve this problem by introducing a stopping time, which is the smallest time at which this condition no longer holds, and prove in Corollary 3 that this stopping time will be $\infty$ with high probability for $\alpha$ sufficiently large. The minimum eigenvalue of the NTK is strictly positive, at least as long as $\omega_{t}-\omega_{0}||_{2}leqr:=\frac{\lambda}{\mathrm{Lip}(Dh)}$. Therefore, we will introduce the random stopping time
@@
denoting the first time $\lambda Id\preceq Dh(\omega_{t})D^{T}h(\omega_{t})$ is degenerate. In addition, the inequality in Assumption 2 also holds at least for $t\leq\tau$. Working with this quantity and proving $\tau=\infty$ with high probability is a major challenge in our analysis. In addition, we define $\min\{T,\tau\}=:\tilde{T}$, where $T>0$ is some arbitrary but fixed runtime of the algorithm. Let $t<\tilde{T}$ in the following.
@@
$\mathbb{E}_{\omega}[\bar{R}(\alpha h(\omega_{t}))]\leq\mathbb{E}_{\omega}[\bar{R}(\alpha h(\omega_{0}))]\exp(-m\lambda^{2}t),$ (3)
@@
As defined, $\omega_{t}$ follows the SDE in equation 1. In the following, we apply Itô’s formula to compute $dg(\omega_{t})$, where $g:\mathbb{R}^{p}\rightarrow\mathbb{R}$ is defined by $g(\omega)=\log(\bar{R}(\alpha h(\omega)))$, for all $t$ with $\bar{R}(\alpha h(\omega_{t}))>0$. The gradient and the Hessian of $g$ are as follows:
$\nabla_{\omega}g(\omega)=$ $\frac{1}{\bar{R}(\alpha h(\omega))}\alpha D^{T}h(\omega)\nabla_{h}R(\alpha h(\omega))$
$\nabla^{2}_{\omega}g(\omega)=$ $-\frac{1}{\bar{R}(\alpha h(\omega))^{2}}\alpha^{2}D^{T}h(\omega)\nabla_{h}R(\alpha h(\omega))\nabla_{h}^{T}R(\alpha h(\omega))Dh(\omega)+\frac{1}{\bar{R}(\alpha h(\omega))}\nabla^{2}_{\omega}R(\alpha h(\omega))$
@@
$dg(\omega_{t})=$ $-\frac{\|D^{T}h(\omega_{t})\nabla_{h}R(\alpha h(\omega_{t}))\|_{2}^{2}}{\bar{R}(\alpha h(\omega_{t}))}dt$ (4)
$+\frac{\eta_{\alpha}}{2}\mathrm{Tr}\left(-\frac{\|\nabla_{h}^{T}R(\alpha h(\omega_{t}))Dh(\omega_{t})\sigma_{\alpha}(\omega_{t})\|_{2}^{2}}{\bar{R}(\alpha h(\omega_{t}))^{2}}\right)dt$ (5)
@@
$+\frac{\sqrt{\eta_{\alpha}}\nabla_{h}^{T}R(\alpha h(\omega_{t}))Dh(\omega_{t})\sigma_{\alpha}(\omega_{t})}{\bar{R}(\alpha h(\omega_{t}))}dW_{t}.$ (7)
@@
$M_{t}:=\sqrt{\eta_{\alpha}}\int_{0}^{t}\frac{\nabla_{h}^{T}R(\alpha h(\omega_{s}))Dh(\omega_{t})\sigma_{\alpha}(\omega_{s})}{\bar{R}(\alpha h(\omega_{s}))}dW_{s},$
@@
$\langle M\rangle_{t}=\eta_{\alpha}\int_{0}^{t}\mathrm{Tr}\left(-\frac{\|\nabla_{h}^{T}R(\alpha h(\omega_{s}))Dh(\omega_{s})\sigma_{\alpha}(\omega_{s}))\|_{2}^{2}}{\bar{R}(\alpha h(\omega_{s}))^{2}}\right)ds,$
@@
$\bar{R}(\alpha h(\omega_{t}))=$ $\bar{R}(\alpha h(\omega_{0}))\exp{\left(-\int_{0}^{t}\frac{\|D^{T}h(\omega_{s})\nabla_{h}R(\alpha h(\omega_{s}))\|_{2}^{2}}{\bar{R}(\alpha h(\omega_{s}))}ds\right)}$
@@
$\int_{0}^{t}\frac{\nabla_{h}^{T}R(\alpha h(\omega_{s}))Dh(\omega_{s})D^{T}h(\omega_{s})\nabla_{h}R(\alpha h(\omega_{s}))}{\bar{R}(\alpha h(\omega_{s}))}ds=\int_{0}^{t}\frac{\|\nabla_{h}R(\alpha h(\omega_{s}))\|^{2}_{Dh(\omega_{s})D^{T}h(\omega_{s})}}{\bar{R}(\alpha h(\omega_{s}))}ds.$
@@
$\int_{0}^{t}\frac{\|\nabla R(\alpha h(\omega_{s}))\|^{2}_{Dh(\omega_{s})D^{T}h(\omega_{s})}}{\bar{R}(\alpha h(\omega_{s}))}ds\geq\lambda^{2}\int_{0}^{t}\frac{\|\nabla R(\alpha h(\omega_{s}))\|^{2}_{2}}{\bar{R}(\alpha h(\omega_{s}))}\geq 2\lambda^{2}m\int_{0}^{t}ds=2\lambda^{2}mt,$
using the Polyak-Lojasiewicz inequality, which holds due to the strong convexity of $R$. By Assumption 2, we get
$\frac{\eta_{\alpha}}{2\alpha^{2}}\int_{0}^{t}\mathrm{Tr}\left(\frac{\sigma_{\alpha}(\omega_{s})^{T}\nabla^{2}_{\omega}R(\alpha h(\omega_{s}))\sigma_{\alpha}(\omega_{s})}{\bar{R}(\alpha h(\omega_{s}))}\right)ds\leq\int_{0}^{t}\lambda^{2}m\ ds$
Consequently we obtain
$\bar{R}(\alpha h(\omega_{t}))$ $\leq\bar{R}(\alpha h(\omega_{0}))\exp{\left(-2\lambda^{2}mt+\lambda^{2}mt+M_{t}-\frac{1}{2}\langle M\rangle_{t}\right)}$
$=\bar{R}(\alpha h(\omega_{0}))\exp{\left(-\lambda^{2}mt+M_{t}-\frac{1}{2}\langle M\rangle_{t}\right)}$
@@
$\mathbb{E}_{\omega}[\bar{R}(\alpha h(\omega_{t})]\xrightarrow[]{t\rightarrow\infty}0,$